The width of a garage is five feet less than twice the length. What is the length of the garage if the area is 700 square feet?
Answer by chana(69)

The width of a garage is five feet less than twice the length. What is the length of the garage if the area is 700 square feet?
sol:
let width = b and length = l
given area, l * b = 700 sq. ft
given b = (2* l) - 5
so on substituting l * ((2* l)-5) = 700
simplifying you get the value of length.